Use two equations in two variables to solve the application.
A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much of each must he use to make 20 liters of a solution that is 40% alcohol? (find how many liters for 50% and 25% solutions).

Answer by Alan3354(69443)

Same problem, different numbers.
----
A merchant wants to mix peanuts worth $3 per pound with jelly beans worth $1.50 per pound to make 30 pounds of a mixture worth $2.10 per pound. How many pounds of each should he use? (find # of lbs used for peanuts and jelly beans).
----------
p + j = 30 --- total mix
3p + 1.5j = 2.1*30 --- total cost
etc.